# COSINE-100 Collaboration

### v2

@@ V Prospects for COSINE-100 and COSINE-100U
The COSINE-100U experiment will be the upgraded version of COSINE-100 Lee et al. (2025). It will be located at the Yemilab underground laboratory, also in South Korea. The new laboratory is 300 m deeper than YangYang, reducing even more the contribution of cosmic rays in the experiment. The main objective of COSINE-100U is to enhance the NaI(Tl) crystals light yield and make all 8 crystals suitable for low energy analysis, reducing the detector threshold, and increasing sensitivity to sub-GeV DM searches.
[…]
@@
Since the expected number of events for the WIMP-electron interaction in NaI(Tl) is much higher for very low energy depositions, as presented in the spectra of Fig. 4, a threshold reduction of the crystals would remarkably enhance the analysis sensitivity. Fig. 9 shows an estimated projection of the constraints for COSINE-100U, assuming a 5 photoelectrons threshold. The prospects indicate a notable improvement in sensitivity for low mass WIMPs, however, it remains uncompetitive with limits established by other experiments.
@@
Figure 9: Sensitivity prospects for the COSINE-100U experiment (red) considering the light mediator model (left) and heavy mediator model (right), for 90% CL. The black data points represent the upper limits set for COSINE-100 in this study. The most stringent constraints from experiments and the models parameters for freeze-in and freeze-out are also shown.
@@
An analysis below the current 0.7 keV COSINE-100 threshold, potentially extended down to a one photoelectron level, is currently being investigated. Given that a precise background modeling at such a small energy deposition is unfeasible due to the very high noise rate, an alternative approach in terms of event selection and data analysis will be required. It is proposed to conduct an annual modulation analysis, which is less reliant on the exact determination of background activity, provided that the stability of noise and background rates are ensured. Once the feasibility of this study is confirmed using COSINE-100 data, it can be implemented in COSINE-100U NaI(Tl) crystals to enhance the analysis sensitivity due to the higher light emission and collection efficiency.
To accurately assess WIMP-electron scattering events that deposit energies below the keV region, it is no longer appropriate to treat sodium and iodine atoms as free within COSINE-100 crystals, as the NaI(Tl) crystal band structure becomes relevant. Studies on the behavior of the NaI(Tl) crystal at low energy depositions have shown promising results of this material when compared to silicon, xenon and argon, which compose the most sensitive experiments to WIMP-electron search Trickle (2022); Derenzo et al. (2017).
@@ VI Conclusion
We reported the search for WIMP scattering on electrons using a 2.82 years data-set of the COSINE-100 experiment NaI(Tl) crystals. We found no excess events consistent with the expected signals for the WIMP-electron interaction assuming a heavy or a light vector boson mediator. 90% confidence level upper limits were set on the WIMP-electron scattering cross section. The most stringent constraints exclude cross sections above 6.4 $\times$ 10-33 cm2 for a 0.25 GeV WIMP, assuming a light mediator, and above 3.4 $\times$ 10-37 cm2 for a 0.4 GeV WIMP, assuming a heavy mediator. These results represent the most restrictive exclusion limits on the WIMP-electron scattering model for a NaI(Tl) target to date, and fully cover the parameters attributed to DAMA/LIBRA modulation results interpreted within this framework. We propose a more sensitive analysis in NaI(Tl) crystals, aiming the investigation of events that generate few photoelectrons in the PMTs. This study could potentially allow the search for MeV WIMPs in NaI(Tl), and will be the subject of a future report.
@@ Acknowledgements.
We thank the Korea Hydro and Nuclear Power (KHNP) Company for providing underground laboratory space at Yangyang and the IBS Research Solution Center (RSC) for providing high performance computing resources. This work is supported by: the Institute for Basic Science (IBS) under project code IBS-R016-A1, NRF-2021R1A2C3010989, NRF-2021R1A2C1013761, RS-2024-00356960, RS-2025-25442707 and RS-2025-16064659, Republic of Korea; NSF Grants No. PHY-1913742, United States; STFC Grant ST/N000277/1 and ST/K001337/1, United Kingdom; Grant No. 2021/06743-1, 2022/12002-7, 2022/13293-5 and 2025/01639-2 FAPESP, CAPES Finance Code 001, CNPq 304658/2023-5, Brazil; LPDP Grant under WCU Equity 2024, Indonesia. *
@@ Appendix A Non-relativistic $f_{ion}$ calculation, and comparison with relativistic model
The analytic calculation method for $f_{ion}$ considered in the non-relativistic spectra presented in Fig. 3 is the same used in the studies reported by DarkSide-50 collaboration Agnes et al. (2018, 2023b). The Roothan-Hartree-Fock (RHF) wavefunctions represent the radial part of bound electrons. These are written as a linear combination of hydrogen-like wavefunctions (Slater-type orbitals) McLean and McLean (1981).
@@
It is possible to insert into the analytic model an effective charge dependent on the distance between the electron and the nucleus, derived from the ”frozen core approximation” Pandey et al. (2020). In this formulation, $Z^{nl}_{eff}(r)$ can be written by:
@@
where $r_{>}$ is the highest value between $r$ and $r^{\prime}$.
@@
Figure 10: Comparison between different methods for performing the $f_{ion}$ calculation for NaI, considering a 2 keV energy deposition. The $f_{ion}$ calculated by the relativistic method Caddell et al. (2023); Roberts and Caddell (2023), by the $Z^{nl}_{eff}$ dependent on $r$ model, and by the fixed $Z^{nl}_{eff}$ model are presented by the blue, orange and green curves, respectively.
Updating the non-relativistic model with a $Z^{nl}_{eff}$ dependent on $r$ is able to make $f_{ion}$ more compatible with the relativistic approach, particularly at intermediate $q$. However, it still lacks a correct modeling of electron wavefunctions near the nucleus, whose contribution is dominant for large $q$. Having in mind that for energy deposition in the keV range $q$ can be up to 50 MeV for $m_{\chi}$ in the GeVs, it is exceptionally important to include relativistic effects in the $f_{ion}$ computation of this COSINE-100 data analysis of the WIMP-electron scattering model.

### v3

@@ V Prospects for COSINE-100 and COSINE-100U
The COSINE-100U experiment will be the upgraded version of COSINE-100 [71]. It will be located at the Yemilab underground laboratory, also in South Korea. The new laboratory is 300 m deeper than YangYang, reducing even more the contribution of cosmic rays in the experiment. The main objective of COSINE-100U is to enhance the NaI(Tl) crystals light yield and make all 8 crystals suitable for low energy analysis, reducing the detector threshold, and increasing sensitivity to sub-GeV DM searches.
A new encapsulation method, focused on removing the need for a quartz light guide between the crystals’ end faces and PMTs, has been proved to reduce scintillation photon losses and improve light collection [72, 73]. The crystals operation at 33 °C-$3\text{\,}\mathrm{\SIUnitSymbolCelsius}$ also provides an increase in light yield [74]. Both upgrades will be applied to COSINE-100U experiment, and are expected to enhance the light yield by about 40%. This is expected to enable a reduction of the analysis threshold from 0.7 keV to approximately 0.2 keV.
@@
Since the expected number of events for the DM-electron interaction in NaI(Tl) is much higher for very low energy depositions, as presented in the spectra of Fig. 5, a threshold reduction of the crystals would remarkably enhance the analysis sensitivity. Fig. 10 shows an estimated projection of the constraints for COSINE-100U, assuming a 5 photoelectrons threshold (approximately 0.2 keV). The prospects indicate a notable improvement in sensitivity for low mass DM, however, it remains uncompetitive with limits established by other experiments.
@@
Figure 10: Sensitivity prospects for the COSINE-100U experiment (red) considering the light mediator model (left) and heavy mediator model (right), for 90% CL. The black data points represent the upper limits set for COSINE-100 in this study. The most stringent constraints from experiments and the models parameters for freeze-in and freeze-out are also shown. The gray shaded region indicates the parameter space expected to be excluded by COSINE-100U. DM particles with parameters above the dotted gray lines are unable to reach Yemilab due to Earth shielding effects (see Appendix B for details). Note that the parameters of the DM velocity distribution and local energy density differ from those used for the dashed line constraints (see the discussion in Sec. III.1).
@@
An analysis below the current 0.7 keV COSINE-100 threshold, potentially extended down to a one photoelectron level, is currently being investigated. Given that a precise background modeling at such a small energy deposition is unfeasible due to the very high noise rate, an alternative approach in terms of event selection and data analysis will be required. It is proposed to conduct an annual modulation analysis, which is less reliant on the exact determination of background activity, provided that the stability of noise and background rates are ensured. Details of the COSINE-100 crystal data below the 8 photoelectron threshold, as well as the modulation analysis, are left for a future publication. Once the feasibility of this study is confirmed using COSINE-100 data, it can be implemented in COSINE-100U NaI(Tl) crystals to enhance the analysis sensitivity due to the higher light emission and collection efficiency.
To accurately model the DM–electron scattering signals down to the a single photoelectron level, the NaI crystal band structure must be taken into account, and treating the Na and I atoms as isolated is no longer a valid approximation, as discussed in Section III.2. Once the information about electronic wave functions of NaI crystal is obtained, open source codes such as QEdark [18] and EXCEED-DM [50] can be used to compute the expected DM–electron energy spectra.
Sensitivity studies on the behavior of NaI at low energy depositions have shown promising prospects of this material compared to silicon, xenon, and argon, which constitute the most sensitive experiments to light DM–electron interactions [75, 49]. The NaI energy band gap is typically calculated to be 5.9 eV [49], lying between that of silicon (1.11 eV [76]) and liquid xenon ($\sim\,$9.22 eV [77]). Particularly for the heavy mediator scenario, if an analysis at a one photoelectron level can be achieved, the most stringent sensitivity is expected to lie in the DM mass range of 10–50 MeV [78], where cross sections below $6\times 10^{-39}$ cm2 remain unconstrained by current experiments.
@@ VI Conclusion
We reported the search for sub-GeV DM scattering on electrons using a 2.82 years data-set of the COSINE-100 experiment NaI(Tl) crystals. We found no excess events consistent with the expected signals for the DM-electron interaction assuming a heavy or a light vector boson mediator. 90% confidence level upper limits were set on the DM-electron scattering cross section. The most stringent constraints exclude cross sections above 6.4 $\times$ 10-33 cm2 for a 0.25 GeV DM, assuming a light mediator, and above 3.4 $\times$ 10-37 cm2 for a 0.4 GeV DM, assuming a heavy mediator. These results represent the most stringent exclusion limits on the DM-electron scattering model for a NaI(Tl) target to date, and fully cover the parameters attributed to DAMA/LIBRA modulation results interpreted within this framework. We propose a more sensitive study using COSINE-100 NaI(Tl) crystals, aimed at searching for annual modulation signals induced by DM electron interactions in few generated photoelectrons in the PMTs. This study could potentially allow the search for MeV DM particles in NaI(Tl), and will be the subject of a future report.
@@ Acknowledgements.
We thank the Korea Hydro and Nuclear Power (KHNP) Company for providing underground laboratory space at Yangyang and the IBS Research Solution Center (RSC) for providing high performance computing resources. We are also very grateful for Ashlee R. Caddell and Benjamin M. Roberts for the helpful discussions and for performing the customized computation of the ionization factors for sodium and iodine. This work is supported by: the Institute for Basic Science (IBS) under project code IBS-R016-A1, NRF-2021R1A2C3010989, NRF-2021R1A2C1013761, RS-2024-00356960, RS-2025-25442707 and RS-2025-16064659, Republic of Korea; NSF Grants No. PHY-1913742, United States; STFC Grant ST/N000277/1 and ST/K001337/1, United Kingdom; Grant No. 2021/06743-1, 2022/12002-7, 2022/13293-5 and 2025/01639-2 FAPESP, CAPES Finance Code 001, CNPq 304658/2023-5, Brazil; LPDP Grant under WCU Equity 2024, Indonesia.
@@ Appendix A Non-relativistic $f_{ion}$ calculation, and comparison with relativistic model
The analytic calculation method for $f_{ion}$ considered in the non-relativistic spectra presented in Fig. 3 is the same used in the studies reported by DarkSide-50 collaboration [48, 26]. The Roothan-Hartree-Fock (RHF) wavefunctions represent the radial part of bound electrons. These are written as a linear combination of hydrogen-like wavefunctions (Slater-type orbitals) [52, 79].
@@
It is possible to insert into the analytic model an effective charge dependent on the distance between the electron and the nucleus, derived from the “frozen core approximation” [80]. In this formulation, $Z^{nl}_{eff}(r)$ can be written by:
@@
where $r$ is the distance between the electron and the nuclear center, $r^{\prime}$ is an integration variable, and $r_{>}$ is the larger of $r$ and $r^{\prime}$.
@@
Figure 11: Comparison between different methods for performing the $f_{ion}$ calculation for NaI, considering a 2 keV energy deposition. The $f_{ion}$ calculated by the relativistic method [47, 51], by the $Z^{nl}_{eff}$ dependent on $r$ model, and by the fixed $Z^{nl}_{eff}$ model are presented by the blue, orange and green curves, respectively.
Updating the non-relativistic model with a $Z^{nl}_{eff}$ dependent on $r$ is able to make $f_{ion}$ more compatible with the relativistic approach, particularly at intermediate $q$. However, it still lacks a correct modeling of electron wavefunctions near the nucleus, whose contribution is dominant for large $q$. Having in mind that for energy deposition in the keV range $q$ can be up to 50 MeV for $m_{\chi}$ in the GeVs, it is exceptionally important to include relativistic effects in the $f_{ion}$ computation of this COSINE-100 data analysis of the DM-electron scattering model.
Appendix B Determination of constraints upper bounds due to Y2L and Yemilab overburden
Before reaching the detectors in underground experiments, DM particles may scatter in the atmosphere and surrounding rock overburden. This Earth shielding effect can attenuate the flux of DM at the detector, limiting the excluded parameter space at larger cross sections. As a result, sub-GeV DM searches commonly report upper bounds on the derived constraints for both DM–electron [67, 23] and DM–nucleon [81, 82] studies. This effect is particularly relevant for DM–electron scattering searches, since for the cross sections accessible to direct detection experiments the impact of Earth shielding cannot be neglected.
The Verne 2.1 code [68, 69, 70] has been used to compute the DM velocity distribution after the particles propagate through the laboratory overburden. A dark photon mediator has been assumed, as well as DM scattering with nucleons, for which the cross section is expected to be significantly larger than ${\overline{\sigma}}_{e}$ for the DM masses considered [83]. The laboratory depths are fixed to 700 m for Y2L and 1000 m for Yemilab. Also, the default Verne rock and atmospheric compositions have been adopted.
Several simplifying assumptions have been adopted in the evaluation of the Earth shielding effect. DM particles have been considered to reach the laboratories at an angle of 54° off the vertical. Although this angle varies during an year and a day, 54° corresponds to the annual average incidence angle. A flat rock overburden has been assumed, neglecting the detailed mountain profiles above Y2L and Yemilab. Finally, the default Verne rock composition has been taken to be uniform along the entire DM trajectory.
To obtain upper bounds on the constraints, the expected energy spectra in the COSINE-100 and COSINE-100U crystals after including the detector response have been generated using the DM velocity distributions calculated with Verne. The upper bound values of the DM–electron cross section are then defined by requiring the total number of events in the ROI to match that obtained for the cross sections that define the constraints in Figs. 9 and 10.
It is important to highlight that Verne 2.1 includes a module designed for calculating the Earth-shielding effect for light DM particles [70], and a separate module appropriate for heavy DM particles [69]. For intermediate DM masses, approximately in the 10 MeV $\sim$ 1 GeV range, which is considered in this work, care must be taken in interpreting the resulting upper bounds, since the approximations employed in both modules may become less accurate. Nevertheless, the upper bounds derived in this analysis lie at least one order of magnitude above the defined constraints and should be interpreted as naive estimates of the Earth shielding effect at Y2L and Yemilab.